OpenStudy (anonymous):

Find the 7th term of the geometric sequence with the given terms. a4 = -8 and a5 = -40; a7 = _______. When I solved this I got -104 but it said it was wrong.

OpenStudy (anonymous):

\[=-200r=\frac{ a5 }{a4 }=\frac{ -40 }{ -8 }=5\] a6=a5*5=-40*5=-200 a7=a6*r=-200*5=-1000

OpenStudy (anonymous):

\[an=a1r ^{n-1}\] \[a4=a1r ^{4-1}=a1r ^{3}=-8\] \[a5=a1r ^{5-1}=a1r ^{4}=-40\] \[\frac{ a1r^4 }{a1r^3 }=\frac{ -40 }{-8 },r=5,\] \[a7=a1r ^{7-1}=a1r^6=a1r^4r^2=-40\left( 5 \right)^2=-40*25=-1000\]

OpenStudy (anonymous):

Ok Ok that makes more sense. I have no idea how I got -104. Thank you for your time!:)